# Neural Graduated Assignment for Maximum Common Edge Subgraphs

### v1

@@ 6.4 More Analysis
In this section, we analyze the hyperparameters to better understand their impact on the performance of NGA. Details of the ablation study are provided in Appendix F, and the analysis of the number of Gumbel samples is shown in Appendix G.3. Here, we summarize the impact of the number of iterations $m$ and the hidden dimensions $d$ in Figure LABEL:fig:_more_analysis.
When $m=0$ , our method excludes the NGA process entirely, relying solely on similarities computed by the graph feature encoder (Eq. (6)). In this scenario, we observe a sharp decrease in performance, highlighting the critical role of NGA in achieving strong results. Unlike traditional GA, which requires dozens of iterations to converge, our method delivers competitive performance with as few as two iterations (layers). This demonstrates that the neuralization of GA with parameterization of temperature is more effective than employing a static parameterization.
Moreover, the hidden dimension $d$ significantly affects the expressiveness of our model. When the hidden dimension $d$ is small, the expressiveness is rather limited. As evidenced in Figure LABEL:fig:NGA_dim when $d=1$, the performance of NGA is comparable to that of traditional GA. However, as $d$ increases, the performance steadily improves until it saturates at higher dimensions. This further validates the effectiveness of the temperature parameterization in NGA in accordance with Proposition 2. Based on this analysis, we set $m=4$ and $d=32$ in all experiments to strike a balance between performance and computational efficiency.
@@ 7 Conclusion
This study introduces Neural Graduated Assignment (NGA), a novel approach to the Maximum Common Edge Subgraph (MCES) problem, of which the scalability remains a challenge for traditional methods. Inspired by the principles of Graduated Assignment (GA), NGA employs a neural architecture that refines the assignment process, effectively adapting GA’s core ideas within a neuralized framework to improve computational efficiency and scalability. Empirical results show that NGA significantly outperforms prior methods in both runtime and accuracy, demonstrating strong performance across tasks like graph similarity and retrieval. The introduction of this versatile method offers substantial contributions to the understanding and resolution of assignment problems, suggesting pathways for future exploration and innovation in related areas.
Limitations.
@@ Appendix A More background
The Graduated Assignment (GA) method is depicted in Section 3, which can be summarized in Alg. 3. One should note that GA is not a neural method and only consists of forward updates.
Algorithm 3 Graduated Assignment
1:  Input: initial assignment $\mathbf{S}$, affinity matrix $\mathbf{A}$, pre-defined annealing coefficients $0<\beta_{0}<...<\beta_{m}$
2:  Output: converged assignment $S$
3:  for $i=0$ to $m$ do
4:     while $\mathbf{S}$ not converge do
5:        $\mathrm{vec}(\mathbf{S})\leftarrow\mathbf{A}\mathrm{vec}(\mathbf{S})$
6:        $\mathbf{S}\leftarrow\exp(\beta_{i}\mathbf{S})$
7:        $\mathbf{S}\leftarrow\mathrm{Sinkhorn}(\mathbf{S})$
8:     end while
9:  end for
@@ Appendix B Association Common Graph Explanation
We aim to show that each valid subgraph $G$ can be transformed into a common subgraph $G^{\prime}$ of $G_{1}$ and $G_{2}$, satisfying $|\mathcal{E}(G)|=|\mathcal{E}(G^{\prime})|$. In this way, finding the MCES of two graphs is reduced to finding the largest subgraph in this set. Let $\mathcal{G}_{1}$ denote the set of all valid subgraphs of $G_{1}\Diamond G_{2}$, and let $\mathcal{G}_{2}$ denote the set of all common subgraphs of $G_{1}$ and $G_{2}$. We claim that there exists a surjective function $f:\mathcal{G}_{1}\to\mathcal{G}_{2}$ that transforms $G$ into $G^{\prime}$, such that $|\mathcal{E}(G)|=|\mathcal{E}(f(G))|$. If this claim holds, each subgraph of the association graph $G_{1}\Diamond G_{2}$ corresponds to a common subgraph of $G_{1}$ and $G_{2}$ without omitting any common subgraph. Consequently, the subgraph with the largest number of edges corresponds to the MCES. Here we give the construction of $f$, and the proof of the surjection of function $f$. The edge set of $G$ is given by =$\mathcal{E}=\{((u_{i},v_{i}),(u_{j},v_{j}))\mid((u_{i},v_{i}),(u_{j},v_{j}))\in\mathcal{E}_{\Diamond}\}$. From this edge set, we restore all the edge information in the original graph $G_{1}$, denoted as =$\mathcal{E}^{\prime}=\{(u_{i},u_{j})\mid((u_{i},v_{i}),(u_{j},v_{j}))\in\mathcal{E}\}\subseteq\mathcal{E}(G_{1}).$ The corresponding vertex set is naturally defined as $\mathcal{V}^{\prime}=\{u_{k}\mid\exists u_{l},(u_{k},u_{l})\in\mathcal{E}^{\prime}\}.$ Thus, the subgraph $G^{\prime}=\{\mathcal{E}^{\prime},\mathcal{V}^{\prime}\}\subseteq G_{1}$ is uniquely determined, and we define $f(G)=G^{\prime}$. Similarly, we can construct $G^{\prime\prime}=\{\mathcal{E}^{\prime\prime}\subseteq\mathcal{E}(G_{2}),\mathcal{V}^{\prime\prime}\subseteq\mathcal{V}(G_{2})\}\subseteq G_{2}$. Since the matching pair in $\mathcal{E}$ is unique, $G^{\prime}$ is isomorphic to $G^{\prime\prime}$. Hence, $G^{\prime}=f(G)$ is the desired common subgraph of $G_{1}$ and $G_{2}$. Furthermore, it is straightforward to verify that $|\mathcal{E}(G)|=|\mathcal{E}(f(G))|$.
@@ C.1 Datasets
The key statistics of datasets used in this paper are summarized in Table 4. The solution time distributions of three datasets are illustrated in Figure LABEL:fig:solve_time.
@@ C.3 Discussion of Baseline Methods
Besides, we provide time complexity comparison between baseline methods and ours in this section. For notation simplicity, we assume that both the input graphs have $N$ nodes and $E$ edges. GA has the same iteration process as our method, which has a complexity of $\mathcal{O}(mN^{4})$. NGM has the same complexity as our method, which is $\mathcal{O}(mkN^{4})$. The rest baseline methods, including RASCAL, FMCS and Mcsplit have a complexity of approximately $\mathcal{O}(2^{N})$ due to the NP-complete nature of the MCES problem. As for our method, the construction of ACG has complexity $\mathcal{O}(N^{2}+E^{2})$. The initial assignments are obtained by GNN with a complexity of $\mathcal{O}(N^{2}d)$. The complexity of NGA iteration is $\mathcal{O}(mkN^{4})$, where $m$ is the number of iterations and $k$ is the number of epochs in training. Overall, the time complexity of our method is $\mathcal{O}(mkN^{4})$. Compared with these baseline methods, our proposed method provides efficient approximations of the solution in polynomial time without relying on exhaustive exploration of the solution space.
@@ D.1 Convergence Behavior under Learnable Parameters
$-\frac{\mathbf{X}^{\top}\mathbf{X}}{2}\rightarrow\min_{\sigma}(-\mathbf{X}^{\top}\sigma+\frac{\sigma^{\top}\sigma}{2})$ (16)
@@
[…]
@@
Step  2:
[…]
ReturntoStep $\mathrm{Return\ to\ Step\ 1\ until\ convergence.}$
@@ D.2 Proof of Faster Convergence for Product Parameterization
Moreover, this proposition is supported by empirical evidence. Specifically, we provide the loss curves in Figure LABEL:fig:empirical_evidence_of_convergence. Compared to parameterizing the temperature as a learnable scalar, our method exhibits faster convergence. Within the same training time, this results in improved performance. As shown in Table 5 from Appendix F, the product parameterization (NGA) outperforms the scalar version (NGA-Scalar) by approximately 15%. These results highlight the effectiveness of our method compared to directly parameterizing $\beta_{l}$ as a learnable scalar.
@@ Appendix F Ablation Study
Finally, to demonstrate the superiority of the product parameterization $\beta_{l}$ in NGA, we introduce two variants: (1) NGA-Manual employs a manually configured temperature schedule initialized with negative values that gradually transitions to positive values, aligned with the mean parameter distribution patterns observed in Fig. LABEL:fig:empirical_evidence. (2) NGA-Scalar parameterizes temperature as a learnable scalar. In Table 5, both variants exhibit suboptimal performance, primarily due to constrained adaptive capacity and limited expressive power. This performance gap underscores the importance of our design. Note that MLP is composed of learnable weights, which is equivalent to our original implementation.
[…]
@@ Appendix G Additional Experimental Results
G.1 Results for Case Study
In this section, we select several instances from the three datasets as case studies. We give a time budget of 3600 seconds and record the best solution found so far for different methods, as shown in Figure LABEL:fig:best_solution_so_far. On the AIDS and MCF-7 datasets, our method can find better solutions within the time budget. On the MOLHIV dataset, both NGA and RASCAL find the best solution, but NGA takes much shorter time.
@@ G.2 Empirical Evidence
As shown in Figure LABEL:fig:empirical_evidence, the parameterized temperature varies with different iterations in NGA, revealing stable trends across the datasets. In the initial iterations of NGA, we observe a cautious exploration of the parameter space, characterized by relatively small and stable values of the parameterized temperature. As the NGA iteration progresses, the algorithm transitions to a more aggressive convergence strategy, reflected by a significant increase in the magnitude of the parameterized temperature. Our empirical results consistently align with the theoretical analysis, highlighting distinct patterns along the optimization trajectory. These patterns reflect the dynamic behavior of the algorithm, where early stages focus on exploration and later stages emphasize convergence, guided by the learned temperature sequence.
G.3 Additional Parameter Analysis

### v2

@@ 6.4 More Analysis
In this section, we first analyze the hyperparameters to better understand their impact on the performance of NGA. Details of the ablation study are provided in Appendix F, and the analysis of the number of Gumbel samples is shown in Appendix G.4. Here, we summarize the impact of the number of iterations $m$ and the hidden dimensions $d$ in Figure 5(a).
(a) How NGA performs on MCES with different parameters.
When $m=0$ , our method excludes the NGA process entirely, relying solely on similarities computed by the graph feature encoder (Eq. (6)). In this scenario, we observe a sharp decrease in performance, highlighting the critical role of NGA in achieving strong results. Unlike traditional GA, which requires dozens of iterations to converge, our method delivers competitive performance with as few as two iterations (layers). This demonstrates that the parameterization of temperature is more effective than employing a static parameterization.
Moreover, the hidden dimension $d$ significantly affects the expressiveness of our model. When the hidden dimension $d$ is small, the expressiveness is rather limited. As evidenced in Figure LABEL:fig:NGA_dim when $d=1$, the performance of NGA is comparable to that of traditional method. However, as $d$ increases, the performance steadily improves until it saturates at higher dimensions. This further validates the effectiveness of the temperature parameterization in NGA in accordance with Proposition 2. Based on this analysis, we set $m=4$ and $d=32$ in all experiments to strike a balance between performance and computational efficiency.
To broaden the scope and potential impact of our work, we have conducted additional experiments on a subset of QAPLIB [8] instances across different categories in Appendix G.5, where our method exhibits competitive performance compared with baseline methods.
@@ 7 Conclusion
This study introduces Neural Graduated Assignment (NGA), a novel approach to the Maximum Common Edge Subgraph (MCES) problem, of which the scalability remains a challenge for traditional methods. Drawing inspirations from annealing mechanism in statistical physics, NGA employs a neural architecture to iteratively refine the assignment process, leveraging high-dimensional learnable parameters to improve computational efficiency, scalability, and adaptive problem-solving. Empirical results show that NGA significantly outperforms prior methods in both runtime and accuracy, demonstrating strong performance across tasks like graph similarity and retrieval. The introduction of this versatile method offers substantial contributions to the understanding and resolution of assignment problems, suggesting pathways for future exploration and innovation in related areas.
@@ Appendix A More background
The Graduated Assignment (GA) is a classical optimization method rooted in principles of statistical physics and commonly applied to assignment problems such as graph matching [18, 9]. GA operates by iteratively updating a soft assignment matrix, which represents the probabilistic relationship between elements of two sets (e.g., nodes, edges) that are being matched. One should note that GA is not a neural method and only consists of forward updates.
@@ Appendix B Association Common Graph Explanation
We aim to show that each valid subgraph $G$ can be transformed into a common subgraph $G^{\prime}$ of $G_{1}$ and $G_{2}$, satisfying $|\mathcal{E}(G)|=|\mathcal{E}(G^{\prime})|$. In this way, finding the MCES of two graphs is reduced to finding the largest subgraph in this set. Let $\mathcal{G}_{1}$ denote the set of all valid subgraphs of $G_{1}\Diamond G_{2}$, and let $\mathcal{G}_{2}$ denote the set of all common subgraphs of $G_{1}$ and $G_{2}$. We claim that there exists a surjective function $f:\mathcal{G}_{1}\to\mathcal{G}_{2}$ that transforms $G$ into $G^{\prime}$, such that $|\mathcal{E}(G)|=|\mathcal{E}(f(G))|$. If this claim holds, each subgraph of the association graph $G_{1}\Diamond G_{2}$ corresponds to a common subgraph of $G_{1}$ and $G_{2}$ without omitting any common subgraph. Consequently, the subgraph with the largest number of edges corresponds to the MCES. Here we give the construction of $f$, and the proof of the surjection of function $f$. The edge set of $G$ is given by $\mathcal{E}=\{((u_{i},v_{i}),(u_{j},v_{j}))\mid((u_{i},v_{i}),(u_{j},v_{j}))\in\mathcal{E}_{\Diamond}\}$. From this edge set, we restore all the edge information in the original graph $G_{1}$, denoted as $\mathcal{E}^{\prime}=\{(u_{i},u_{j})\mid((u_{i},v_{i}),(u_{j},v_{j}))\in\mathcal{E}\}\subseteq\mathcal{E}(G_{1}).$ The corresponding vertex set is naturally defined as $\mathcal{V}^{\prime}=\{u_{k}\mid\exists u_{l},(u_{k},u_{l})\in\mathcal{E}^{\prime}\}.$ Thus, the subgraph $G^{\prime}=\{\mathcal{E}^{\prime},\mathcal{V}^{\prime}\}\subseteq G_{1}$ is uniquely determined, and we define $f(G)=G^{\prime}$. Similarly, we can construct $G^{\prime\prime}=\{\mathcal{E}^{\prime\prime}\subseteq\mathcal{E}(G_{2}),\mathcal{V}^{\prime\prime}\subseteq\mathcal{V}(G_{2})\}\subseteq G_{2}$. Since the matching pair in $\mathcal{E}$ is unique, $G^{\prime}$ is isomorphic to $G^{\prime\prime}$. Hence, $G^{\prime}=f(G)$ is the desired common subgraph of $G_{1}$ and $G_{2}$. Furthermore, it is straightforward to verify that $|\mathcal{E}(G)|=|\mathcal{E}(f(G))|$.
@@ C.1 Datasets
The key statistics of datasets used in this paper are summarized in Table 4. The solution time distributions of three datasets are illustrated in Figure 9(a).
(a) Solution time distributions on three datasets
@@ C.3 Discussion of Baseline Methods
Besides, we provide time complexity comparison between baseline methods and ours in this section. For notation simplicity, we assume that both the input graphs have $N$ nodes and $E$ edges. For molecular graphs where atoms typically have only a few neighbors (e.g., due to chemical valency constraints), the constructed ACG is relatively sparse, and we denote its number of edges as $M$. The computation involving the ACG therefore has a complexity of $\mathcal{O}(M)$. The initial assignments are obtained by GNN with a complexity of $\mathcal{O}(N^{2}d)$. Therefore, the overall complexity is $\mathcal{O}(M+N^{2})$. The complexity of NGA iteration is $\mathcal{O}(mk(M+N^{2}))$, where $m$ is the number of iterations and $k$ is the number of epochs in training. Overall, the time complexity of our method is $\mathcal{O}(mk(M+N^{2}))$. GA has the same iteration process as our method, which has a complexity of $\mathcal{O}(m(M+N^{2}))$. NGM has the same complexity as our method, which is $\mathcal{O}(mk(M+N^{2}))$. The rest baseline methods, including RASCAL, FMCS and Mcsplit have a complexity of approximately $\mathcal{O}(2^{N})$ due to the NP-complete nature of the MCES problem. Compared with these baseline methods, our proposed method provides efficient approximations of the solution in polynomial time without relying on exhaustive exploration of the solution space.
@@ D.1 Convergence Behavior under Learnable Parameters
$\frac{\mathbf{X}^{\top}\mathbf{X}}{2}\rightarrow\min_{\sigma}(-\mathbf{X}^{\top}\sigma+\frac{\sigma^{\top}\sigma}{2})$ (16)
@@
We use $deg_{i}$ to represent the degree of node $i$ and $deg\textsubscript{max}=\max_{i\in\mathcal{V}_{\Diamond}}deg_{i}$. By setting $\gamma=deg\textsubscript{max}+\epsilon$, where $\epsilon>0$ is a small quantity, the following equation holds for any nonzero vector $\mathbf{x}$:
@@
$\textbf{Step \ 1: }\sigma\leftarrow\mathrm{vec}(\mathbf{S})$ (24)
Step 2:
$\quad\textbf{Step \ 2a: }\hat{\mathbf{S}}_{ai}\leftarrow\mathrm{exp}(\beta_{l}\sum_{bj}(\mathcal{A}_{\Diamond}^{(\gamma)})_{ai,bj}\mathbf{\sigma}_{bj})$
@@
$\mathrm{Return\ to\ Step\ 1\ until\ convergence.}$
@@ D.2 Proof of Faster Convergence for Product Parameterization
Moreover, this proposition is supported by empirical evidence. Specifically, we provide the loss curves in Figure 16(b). Compared to parameterizing the temperature as a learnable scalar, our method exhibits faster convergence. Within the same training time, this results in improved performance. As shown in Table 5 from Appendix F, the product parameterization (NGA) outperforms the scalar version (NGA-Scalar) by approximately 15%. These results highlight the effectiveness of our method compared to directly parameterizing $\beta_{l}$ as a learnable scalar.
@@ Appendix F Ablation Study
To demonstrate the effectiveness of the exploration and exploitation process of our proposed NGA, we tested a variant of our method where we use the Sigmoid function to restrict $\beta_{l}>0$, denoted as NGA-Positive. This variant disables the exploration ability of our method. We observe that restricting $\beta_{l}$ to positive values significantly degrades performance, indicating that allowing negative temperatures is crucial for effective exploration and for enabling NGA to escape local minima.
@@
Finally, to demonstrate the superiority of the product parameterization $\beta_{l}$ in NGA, we introduce two variants: (1) NGA-Manual employs a manually configured temperature schedule initialized with negative values that gradually transitions to positive values, aligned with the mean parameter distribution patterns observed in Fig. 16(a). (2) NGA-Scalar parameterizes temperature as a learnable scalar. In Table 5, both variants exhibit suboptimal performance, primarily due to constrained adaptive capacity and limited expressive power. This performance gap underscores the importance of our design. Note that MLP is composed of learnable weights, which is equivalent to our original implementation.
Table 5: Results of Ablation Study in terms of Accuracy $(\%)$ $\uparrow$ for solving MCES problem.
@@ Appendix G Additional Experimental Results
G.1 Additional Results for Solving MCES
We have additionally conducted a detailed analysis of the quality of the solutions obtained by our proposed NGA. In particular, we examine how frequently the MCES identified by NGA coincides with the true optimal solution. This evaluation provides deeper insights into the effectiveness and reliability of NGA beyond average performance metrics. The results highlights not only the overall accuracy of the method but also its stability across different problem instances. A summary of these results is provided in Table 6.
Table 6: The percentage of optimal solutions of MCES obtained on different datasets.
AIDS MCF-7 MOLHIV
49% 56% 61%
G.2 Results for Case Study
In this section, we select several instances from the three datasets as case studies. We give a time budget of 3600 seconds and record the best solution found so far for different methods, as shown in Figure 12(a). On the AIDS and MCF-7 datasets, our method can find better solutions within the time budget. On the MOLHIV dataset, both NGA and RASCAL find the best solution, but NGA takes much shorter time.
@@ G.2 Results for Case Study
(a) Comparison of the best MCES solution size found so far on different dataset.
@@ G.3 Empirical Evidence
As shown in Figure 16(a), the parameterized temperature varies with different iterations in NGA, revealing stable trends across the datasets. In the initial iterations of NGA, we observe a cautious exploration of the parameter space, characterized by relatively small and stable values of the parameterized temperature. As the NGA iteration progresses, the algorithm transitions to a more aggressive convergence strategy, reflected by a significant increase in the magnitude of the parameterized temperature. Our empirical results consistently align with the theoretical analysis, highlighting distinct patterns along the optimization trajectory. These patterns reflect the dynamic behavior of the algorithm, where early stages focus on exploration and later stages emphasize convergence, guided by the learned temperature sequence.
(a) Distribution of parameterized temperature values across iterations in NGA on different datasets.
(b) Loss curves across epochs on different datasets. NGA-Scalar represents an variant of NGA by setting the temperature value as a learnable scalar.
G.4 Additional Parameter Analysis
@@
G.5 Results on QAP Instances
To demonstrate the applicability of our method to address the general QAP problem, we conducted additional experiments on Koopmans-Beckmann’s QAP benchmarks: QAP20, QAP50, and QAP100, corresponding to problems with 20, 50, and 100 nodes. In each task, location node coordinates are uniformly sampled from the unit square $0,1]^{2}$. The flow $f_{ij}$ between facility $i$ and $j$ is sampled uniformly from $[0,1]$, symmetrized, diagonal entries set to zero, and randomly set to zero with probability $p=0.7$. We train with up to 5120 instances and evaluate on a test set of 256 instances from the same distribution. Since these tasks do not involve the MCES-specific structure, we omit ACG construction and directly use the NGA module (Alg. 1). We compare our method with the heuristic solver SM [26] and the learning-based solver NeuOpt [29]. The results are summarized in Table 7. Our method achieves competitive or superior solutions across these QAP instances.
Table 7: Total Assignment Cost (lower is better) on Koopmans-Beckmann’s QAP benchmarks
Method QAP20 QAP50 QAP100
SM 70.07 446.62 1800.75
NeuOpt 61.37 430.61 1789.22
NGA 62.17 407.59 1726.54
Besides, we conduct additional experiments on a subset of QAPLIB instances across different categories. The compared methods include GLAG [15], PATH [48], FAQ [44]. The results are summarized in Table 8. These results show that while our method is primarily designed for MCES, it can still perform competitively on general QAP, further demonstrating the versatility of our approach.
Table 8: Total Assignment Cost (lower is better) on 16 Benchmark Examples of the QAPLIB Library
QAP PATH GLAG FAQ NGA
chr12c 18048 61430 13088 17598
chr15a 19086 78296 29018 17352
chr15c 16206 82452 11936 20978
chr20b 5560 13728 2764 3366
chr22b 8500 21970 8774 8580
esc16b 300 320 314 292
rou12 256320 353998 254336 251094
rou15 391270 521882 371458 389048
rou20 778284 1019622 759838 813576
tai10a 152534 218604 157954 145270
tai15a 419224 544304 397376 418618
tai17a 530978 708754 520754 537896
tai20a 753712 1015832 736140 802464
tai30a 1903872 2329604 1908814 1861206
tai35a 2555110 3083180 2531558 2789834
tai40a 3281830 4001224 3237014 3214724
Appendix H Limitations